OpenStudy (anonymous):

Write a word problem describing the time it takes to complete an activity individually and with a friend.

OpenStudy (anonymous):

Write a word problem describing the time it takes to complete an activity individually and with a friend. For example, if John takes 2 hours to mow his lawn and it takes his sister Maria 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together? Write a rational equation based upon the word problem you created. The rational equation based upon the scenario above would be one half plus one fourth equals one over x.. Each fraction represents the amount of the lawn mowed in one hour. The fraction of one half is John’s portion. Maria’s share is represented by one fourth. The time it would take for both of them to mow the lawn is represented by one over x. Solve the rational equation. Show your work.

OpenStudy (anonymous):

Here is a example please don't copy: It takes a novice painter to paint a whole house in 8 hours when, it takes a professional painter paints a house alone it gets done in 6 hours. How long would it take for a novice and professional painter a house together?

OpenStudy (anonymous):

1/8+1/6=1/x x=24/7 so that's how long it would take...right?

OpenStudy (anonymous):

Yes correct but please don't copy

OpenStudy (anonymous):

If you like what I gave you can you give me a medal please?

OpenStudy (anonymous):

okay thanks dude :D

OpenStudy (anonymous):

Anytime